Solve this Diffrential equation (1-x^2)y"-2xy'+n(n+1)y=0 using Power Series Method


1
Expert's answer
2021-12-24T05:09:53-0500

Replacing 'n' given in equation with 'a' and solving.

"\left(1-x^{2}\right) y^{\prime \prime}-2 x y^{\prime}+b y=0 \text { with } b=a(a+1)"

When x is small, this is "y^{\prime \prime}+b y=0" which has solutions "y=u \sin (x \sqrt{b})+v \cos (x \sqrt{b})" . So, this seems like what the solutions look like.


Proceeding mechanically, let "y(x)=\sum_{n=0}^{\infty} a_{n} x^{n}=a_{0}+a_{1} x+\sum_{n=2}^{\infty} a_{n} x^{n}" . The reason for this will appear later.

"\begin{aligned}

x y^{\prime}(x) &=x \sum_{n=1}^{\infty} n a_{n} x^{n-1} \\

&=\sum_{n=1}^{\infty} n a_{n} x^{n} \\

&=a_{1} x+\sum_{n=2}^{\infty} n a_{n} x^{n} \\

&=\left(1-x^{2}\right) \sum_{n=2}^{\infty} n(n-1) a_{n} x^{n-2} \\

&=\sum_{n=2}^{\infty} n(n-1) a_{n} x^{n-2}-\sum_{n-2}^{\infty} n(n-1) a_{n} x^{n} \\

&=\sum_{n-0}^{\infty}(n+2)(n+1) a_{n+2} x^{n}-\sum_{n-2}^{\infty} n(n-1) a_{n} x^{n} \\

&=2 a_{2}+6 a_{3} x+\sum_{n=2}^{\infty}\left(n(n-1) a_{n}-(n+2)(n+1) a_{n+2}\right) x^{n}

\end{aligned}"


Adding these up, and separating the terms for n<2,

"\begin{gathered}

0=\left(1-x^{2}\right) y^{\prime \prime}-2 x y^{\prime}+b y \\

=2 a_{2}+6 a_{3} x-2 a_{1} x+b\left(2 a_{0}+6 a_{1} x\right)+\sum_{n=2}^{\infty}\left(n(n-1) a_{n}-(n+2)(n+1) a_{n+2}-2 n a_{n}\right. \\

\left.+b a_{n}\right) x^{n} \\

=2\left(a_{2}+b a_{0}\right)+2 x\left(6 a_{3}-2 a_{1}+6 b a_{1}\right)+ \\

\sum_{n=2}^{\infty}\left((n(n-1)-2 n+b) a_{n}-(n+2)(n+1) a_{n+2}\right) x^{n} \\

=2\left(a_{2}+b a_{0}\right)+2 x\left(6 a_{3}-2 a_{1}(1-3 b)\right)+\sum_{n=2}^{\infty}\left((n(n-3)+b) a_{n}-(n+2)(n+1) a_{n+2}\right) x^{n}

\end{gathered}"

Equating the coefficients to zero,

"\begin{aligned}
&a_{2}+b a_{0}=0 \text { or } a_{2}=-b a_{0} ; \\
&0=6 a_{3}-2 a_{1}+6 b a_{1}=6 a_{3}-2 a_{1}(1-3 b) \text { or } a_{3}=a_{1}\left(\frac{1}{3}-b\right) ; \\
&\text { and, for } n \geq 2,0=(n(n-3)+b) a_{n}-(n+2)(n+1) a_{n+2} \text { or } a_{n+2}=\frac{n(n-3)+b}{(n+2)(n+1)} a_{n} . \\
&\text { From this, we see that there are two independent solutions: One with } a_{0}=0 \text { (the odd solution) } \\
&\text { and one with } a_{1}=0 \text { (the even solution). } \\
&\text { We also see that if } b=\frac{1}{3}, \text { the odd solution is just } a_{1} x / 3, \text { since all the higher odd coefficients are } \\
&\text { zero. } \\
&\text { Similarly, if } b=-n(n-3) \text { for some } n, \text { the even solution is a polynomial, since } a_{n+2}=0 \text { for the } n \\
&\text { such that } b=n(n-3) \text { and all greater even } n .
\end{aligned}"